1. The marginal cost of a trader has been found to be MC = 3Q2 +8Q+ 400 . Determine the total variable cost of producing 100 units of the trade’s product.

2. If the fixed cost of manufacturing a product is ETB 10, 000 and the marginal cost at Q units of output is ETB(60+2.5Q). Find:

a. The function for the total cost of manufacturing x units.

b. The total cost of 200 units.

3. Imagine a perfectly competitive firm producing good A with cost function TC=400+20Q-2Q 2 +2/3Q 3 , where Q is quantity produced

a. determine the firm’s short run supply curve

b. What is the profit maximizing level of output when price of A is birr 180?



1
Expert's answer
2021-12-22T06:08:24-0500

1.Integrating the function:

"Q^3+4Q^2+400Q=100^3+4\times100^2+400\times100=1 080 0000"


2.a. The function for the total cost of manufacturing x units.

"TC=\intop60+2.5Q"

"TC=1.25Q^2+60Q"


b. The total cost of 200 units.  

"=1.25(200)^2+60(200)"

"=62,000"


3.a. the firm’s short run supply curve is MC

"MC=TC'=(400+20Q-2Q2+2\/3Q3)'=20-4Q+2Q^2"


b.

MC=P

"20-4Q+2Q^2=180"

"20-4Q+2Q^2-180=0"

"2Q^2-4Q-160=0"

"Q^2-2Q-80=0"

Q=10

"profit=MR-MC=180\times10-(20-4\times10+2(10^2))=1800-20+40-200=1620"